OpenStudy (anonymous):

Check my answer please :) Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226. -334 -274 -284 --> -346 my answer

OpenStudy (***[isuru]***):

ok... we have common difference as (-226 - 26)/21 = -12 and the first term is 26 so.. tn = a + (n-1)d t31 = 26 + (30)(-12) = -334 so... u need to recheck ur answer :)

OpenStudy (anonymous):

ohh, okay but why is a1 negative?

OpenStudy (***[isuru]***):

huh... between a1 and a22 the common difference is repeated 21 times... so if u subtract a1 from a22 and divide it by 21 .. u get the common difference :)

OpenStudy (anonymous):

okay thank you :)

OpenStudy (***[isuru]***):

u r welcome